Thomas Taylor is planning to invest $29,000 today in a mutual fund that will provide a return of 8 percent each year. What will be the value of the investment in 10 years? (If you solve this problem with algebra round intermediate calculations to 6 decimal places, in all cases round your final answer to the nearest penny.)
Answer #1
Future value FV= PV × (1+r)^N
Where as:
Present value PV=                         29,000
Rate of interest r= 8.00%
Number of years N=                           10.00
Future value FV= 29000 × (1+0.08)^10
FV=                   62,608.82

Answer is $62,608.82

please rate.
Answer #2

SOLUTION :


Value in 10 years 

= A(1+r)^10

= 29000*(1 + 0.08)^10

= 62608.82 ($) (ANSWER)

answered by: Tulsiram Garg